OpenStudy (anonymous):

guys help me !!! 2x+3y=9 3x+2y=9

OpenStudy (saifoo.khan):

0

myininaya (myininaya):

\[3y=9-2x=> 3y=-2x+9=> y=\frac{-2}{3}x+3\] \[2y=9-3x =>y=\frac{-3}{2}x+\frac{9}{2}\] set them equal and solve for x \[\frac{-2}{3}x+3=\frac{-3}{2}x+\frac{9}{2}\] \[\frac{-2}{3}x+\frac{3}{2}x=\frac{9}{2}-3\] \[\frac{-4+9}{6}x=\frac{9-6}{2}\] \[\frac{5}{6}x=\frac{3}{2} => x=\frac{3}{2} \cdot \frac{6}{5}=\frac{6}{2} \cdot \frac{3}{5}=3 \cdot \frac{3}{5}=\frac{9}{5}\] now you can find y by plugin' the x we just found

myininaya (myininaya):

\[y=\frac{-2}{3}(\frac{9}{5})+3\]